OpenStudy (anonymous):

How would I go about solving this? (7/3)-(3/x-5)=(19/12)

OpenStudy (nathan917):

Thats easy! :P Step by Step?

OpenStudy (nathan917):

Or all at once?

OpenStudy (nathan917):

And welcome to open study!

OpenStudy (anonymous):

i just want to know what i multiply. I know i need a common denominator but when i multiply what all do i multiply. And thanks :)

OpenStudy (nathan917):

Multiply through by x-5, remembering that x=5 can never be a solution to this equation from the start: 7(x-5)/3 - 3 = 19(x-5)/12 Multiply through by 12: 28(x-5) - 36 = 19(x-5) Expand: 28x - 140 - 36 = 19x - 95 Collect like terms: 9x = 81 x = 9

OpenStudy (nathan917):

This a good way to explain it?

OpenStudy (nathan917):

12 in the denominator.

OpenStudy (nathan917):

So multiply everything by 12

OpenStudy (anonymous):

Alright i think i get it. My teacher isn't the best at explaining so i normally teach myself this stuff. so thanks so much :)

OpenStudy (nathan917):

Anytime.